Meera is now 10 years older than her twin brothers.After 3 years from now,thrice her age will be four times the sum of her brothers age then.find meera's present age
Answer by ankor@dixie-net.com(22740)

Meera is now 10 years older than her twin brothers.
m = b + 10
After 3 years from now, thrice her age will be four times the sum of her brothers age then.
3(m + 3) = 4(2(b+3))
3m + 9 = 8b + 24
replace m with (b+10)
3(b+10) + 9 = 8b + 24
3b + 30 + 9 = 8b + 24
39 - 24 = 8b - 3b
15 = 5b
b = 15/5
b = 3 yrs is the twin brothers' age
then
10 + 3 = 13 is Meera's age
:
:
Check this, in 3 yrs, m is 16 and b is 6
3(16) = 4(6+6)
48 = 48